# Language Model Council: Benchmarking Foundation Models on Highly Subjective Tasks by Consensus

### v1

@@ Acknowledgements
Justin led the research. Amanda and Flor led the human study. We thank Sahand Sabour for creating EmoBench and insightful discussions about emotionally rich synthetic data. We thank Alex Tamkin for proposing the idea to measure the value of the incremental judge. We thank Mitchell Gordon for his suggestion to qualitatively analyze why certain responses are preferred. We thank David So for his ideas on calibration, repeatability, and model affinity. We thank Federico Bianchi and Sam Paech for their idea of measuring oligarchical councils and for their feedback on prompt design. Sam Paech provided insightful discussions on separability, voting aggregation, length bias, and the relationship to other leaderboards, and he advised us throughout the project. Finally, we thank Predibase for their support and funding this research.
Appendix A Language Model Council Configurations
@@
United States Open AI gpt-4o-2024-05-13 05/24 1287 88.7 [Uncaptioned image] Proprietary
United States Open AI gpt-4-turbo-04-09 04/24 1256 [Uncaptioned image] [Uncaptioned image] Proprietary
United States Open AI gpt-4-0613 06/23 1246 86.4 [Uncaptioned image] Proprietary
United States Open AI gpt-3.5-turbo-0125 01/24 1102 70.0 [Uncaptioned image] Proprietary
France Mistral mistral-large-latest 02/24 1156 81.2 [Uncaptioned image] Proprietary
France Mistral open-mixtral-8x22b 04/24 1146 77.8 176 B Apache 2.0
France Mistral open-mixtral-8x7b 12/23 1114 70.6 56 B Apache 2.0
United States Meta llama-3-70b-chat-hf 04/24 1208 82.0 70 B Llama 3 Community
United States Meta llama-3-8b-chat-hf 04/24 1153 68.4 8 B Llama 3 Community
United States Google gemini-1.5-pro-preview-0409 05/24 1268 81.9 [Uncaptioned image] Proprietary
United States Google gemini-1.0-pro 04/24 1208 71.8 [Uncaptioned image] Proprietary
United States Databricks dbrx 03/24 1103 73.7 132 B DBRX LICENSE
Canada Cohere command-r-plus 04/24 1189 75.7 104 B CC-BY-NC-4.0
Canada Cohere command-r 04/24 1147 68.2 35 B CC-BY-NC-4.0
United States Anthropic claude-3-opus-20240229 03/24 1248 86.8 [Uncaptioned image] Proprietary
United States Anthropic claude-3-sonnet-20240229 03/24 1201 79.0 [Uncaptioned image] Proprietary
United States Anthropic claude-3-haiku-20240307 03/24 1178 75.2 [Uncaptioned image] Proprietary
China Alibaba qwen1.5-110B-chat 02/24 1164 80.2 100 B Qianwen LICENSE
China Alibaba qwen1.5-72B-chat 02/24 1152 77.4 72 B Qianwen LICENSE
China Alibaba qwen1.5-32B-chat 02/24 1126 74.3 32 B Qianwen LICENSE
@@
Table 9: Additional council variations consisting of a hand-picked subset of LLMs.
Appendix B LLM Judge Calibration
To understand the consistency and variability of LLM model judges under different temperatures and pairwise comparison options, we collect pairwise preference ratings on three responses to the same interpersonal conflict. Two responses are competitive, and one is intentionally generic to serve as a ranking baseline (Figure 6). We measure:
Invariability:
How reliably does the model give the same answer every time with the same pair in the same order?
Let:
  • $P$ be the set of all pairs of responses.
  • $n$ be the number of repetitions.
The frequency of the mode for the $i$-th pair is given by:
$f_{i}=\frac{\sum_{j=1}^{n}\mathbb{I}(R_{i,j}=\text{mode}(R_{i}))}{n}$
$invariability=\frac{1}{|P|}\sum_{i\in P}f_{i}$
Consistency:
Does the model give a consistent answer when the order of the same pair of respondents is flipped?
To measure consistency over several repetitions of the same items in both orders, we take the percentage of consistent couplets over all possible rating couplets.
Let:
  • $P$ be the set of all pairs of responses.
  • $n$ be the number of repetitions.
[…]
$consistency=\frac{1}{|P|\cdot n^{2}}\sum_{i\in P}\sum_{j=1}^{n}\sum_{k=1}^{n}\text{are\_consistent}(R_{i,j},R_{i^{\prime},k})$
To assess invariability, each LLM judge is prompted 5 times with the original pairwise comparison prompt (Figure 24 and 5 times with a trivially reworded version of the prompt  [^5] Trivial rewording involves changing the first sentence of the judging prompt (Figure 24) to: ”This person is experiencing an emotional dilemma and is seeking guidance and help.”. To assess consistency, each pair of responses is evaluated in both orders, also for 10 repetitions for each.
In our calibration, there are 10 reps of each pair of responses in both positions, so there are $10*10=100$ instances of swapped-position rating couplets. The are_consistent function for consistency metrics is based on the mapping defined in Table 10.
A>>B or A>B A>>B or A>B FALSE TRUE TRUE FALSE
A>>B or A>B B>>A or B>A TRUE FALSE FALSE FALSE
A>>B or A>B Asimilar-$\sim$=B FALSE TRUE TRUE FALSE
B>>A or B>A A>>B or A>B TRUE FALSE FALSE FALSE
B>>A or B>A B>>A or B>A FALSE TRUE FALSE TRUE
B>>A or B>A Asimilar-$\sim$=B FALSE TRUE FALSE TRUE
Asimilar-$\sim$=B A>>B or A>B FALSE TRUE TRUE FALSE
Asimilar-$\sim$=B B>>A or B>A FALSE TRUE FALSE TRUE
Asimilar-$\sim$=B Asimilar-$\sim$=B TRUE FALSE FALSE FALSE
Table 10: Reference table for categorizing a couplet of order-swapped ratings of the same set of items, (A, B) vs. (B, A). We do not penalize consistency as long as the overall side of the preference is consistent. All inconsistent votes are either biased towards the first or second position.
  • Coarse preferences with tie option (A>B, B>A, Asimilar-$\sim$=B)
  • Coarse preferences without tie option (A>B, B>A)
  • Granular preferences without tie option (A>>B, A>B, B>A, B>>A)
Figure 7: Calibration scores for invariability (left) and consistency (right), averaged over 20 LLMs and 10 repetitions for each under different pairwise comparison options.
Somewhat surprisingly, several models do not produce fully invariant ratings, even with $temperature=0$. However, using a lower temperature increases consistency. Excluding the tie option slightly improves invariability and consistency at some temperatures, with no negative impact at $temperature=0$. Table/Figure [table ref] lists detailed calibration results for individual LLMs with granular pairwise comparison options without ties and with $temperature=0$.
Our calibration study concludes with the decision to use granular comparison options without a tie to "force" judges to choose a side, thereby better distinguishing models, and with $temperature=0$.
(strong votes) (first) (second)
claude-3-haiku 100.0% 50.0% 50.0% 50.0% 50.0%
claude-3-opus 100.0% 50.0% 100.0% 0.0% 0.0%
claude-3-sonnet 100.0% 50.0% 100.0% 0.0% 0.0%
command-r 100.0% 50.0% 50.0% 50.0% 50.0%
command-r-plus 100.0% 50.0% 100.0% 0.0% 0.0%
mistral-large 100.0% 50.0% 50.0% 0.0% 0.0%
mistral-medium 100.0% 50.0% 50.0% 0.0% 0.0%
mixtral-8x7b 100.0% 25.0% 50.0% 0.0% 0.0%
gpt-3.5-turbo-0125 82.5% 50.0% 95.0% 0.0% 0.0%
gpt-4-0613 100.0% 50.0% 100.0% 0.0% 0.0%
gpt-4-turbo-2024-04-09 92.5% 50.0% 100.0% 0.0% 0.0%
gpt-4o-2024-05-13 95.0% 50.0% 100.0% 0.0% 0.0%
qwen1.5-110B-Chat 100.0% 50.0% 100.0% 0.0% 0.0%
qwen1.5-32B-Chat 95.0% 25.0% 100.0% 0.0% 0.0%
qwen1.5-72B-Chat 100.0% 50.0% 50.0% 0.0% 0.0%
dbrx-instruct 92.5% 50.0% 65.0% 50.0% 50.0%
llama-3-70b-chat 100.0% 50.0% 100.0% 0.0% 0.0%
llama-3-8b-chat 82.5% 50.0% 50.0% 50.0% 50.0%
gemini-1.0-pro 75.0% 25.0% 69.5% 0.0% 0.0%
gemini-1.5-pro 100.0% 50.0% 100.0% 0.0% 0.0%
Table 11: Judging calibration results for 20 LLMs with using granular comparison options without a tie, with $temperature=0$.
Appendix C Further Details on Main Experiment
@@
Table 12: Extended Council judging profiles for bias, with and without consistent votes.
@@
Table 13: Extended Council judging profiles for agreement, with and without consistent votes.
@@
Table 14: Extended Council judging profiles for affinity, with and without consistent votes.
Figure 9: Full heatmap of LLM judge to LLM respondent affinities from the main experiment.
Figure 10: Full heatmap of LLM judge to LLM respondent affinities from the main experiment, with Council consensus affinity subtracted out.
Figure 11: Graph of top 5 affinities. An edge exists from LLM $a$ to LLM $b$ if $\text{affinity}(a,b)$ is in the top 5 affinities for LLM $a$.
Figure 12: Full heatmap of LLM judge to LLM judge Cohen’s Kappa sidewise agreement scores from the main experiment.
Figure 13: Graph of top 5 agreement. An edge exists from LLM $a$ to LLM $b$ if $\text{agreement}(a,b)$ is in the top 5 agreement for LLM $a$.
Figure 14: Expected LLM respondent vs. LLM respondent win rates derived from Terry-Bradley coefficients.
Appendix D Human Evaluation

### v2

@@ Acknowledgements
Justin led the research. Amanda and Flor led the human study. Benji ran Monte Carlo simulaitons and generated gradient charts. We thank Sahand Sabour for creating EmoBench and insightful discussions about emotionally rich synthetic data. We thank Alex Tamkin for proposing the idea to measure the value of the incremental judge. We thank Mitchell Gordon for his suggestion to qualitatively analyze why certain responses are preferred. We thank David So for his ideas on calibration, repeatability, and model affinity. We thank Federico Bianchi and Sam Paech for their idea of measuring oligarchical councils and for their feedback on prompt design. Sam Paech provided insightful discussions on separability, voting aggregation, length bias, and the relationship to other leaderboards, and he advised us throughout the project. Finally, we thank Predibase for their support and funding this research.
Appendix A Language Model Council Membership
@@
United States Open AI gpt-4o-2024-05-13 OpenAI (2024b) 05/24 1287 88.7 [Uncaptioned image] Proprietary
United States Open AI gpt-4-turbo-04-09 OpenAI (2024a) 04/24 1256 [Uncaptioned image] [Uncaptioned image] Proprietary
United States Open AI gpt-4-0613 OpenAI et al. (2023) 06/23 1246 86.4 [Uncaptioned image] Proprietary
United States Open AI gpt-3.5-turbo-0125 OpenAI (2023) 01/24 1102 70.0 [Uncaptioned image] Proprietary
France Mistral mistral-large-latest AI (2024) 02/24 1156 81.2 [Uncaptioned image] Proprietary
France Mistral open-mixtral-8x22b Mistral (2024) 04/24 1146 77.8 176 B Apache 2.0
France Mistral open-mixtral-8x7b Jiang et al. (2024) 12/23 1114 70.6 56 B Apache 2.0
United States Meta llama-3-70b-chat-hf Dubey et al. (2024) 04/24 1208 82.0 70 B Llama 3 Community
United States Meta llama-3-8b-chat-hf Dubey et al. (2024) 04/24 1153 68.4 8 B Llama 3 Community
United States Google gemini-1.5-pro-preview-0409 Google (2024b) 05/24 1268 81.9 [Uncaptioned image] Proprietary
United States Google gemini-1.0-pro Google (2024a) 04/24 1208 71.8 [Uncaptioned image] Proprietary
United States Databricks dbrx Databricks (2024) 03/24 1103 73.7 132 B DBRX LICENSE
Canada Cohere command-r-plus Cohere (2024b) 04/24 1189 75.7 104 B CC-BY-NC-4.0
Canada Cohere command-r Cohere (2024a) 04/24 1147 68.2 35 B CC-BY-NC-4.0
United States Anthropic claude-3-opus-20240229 Anthropic (2024) 03/24 1248 86.8 [Uncaptioned image] Proprietary
United States Anthropic claude-3-sonnet-20240229 Anthropic (2024) 03/24 1201 79.0 [Uncaptioned image] Proprietary
United States Anthropic claude-3-haiku-20240307 Anthropic (2024) 03/24 1178 75.2 [Uncaptioned image] Proprietary
China Alibaba qwen1.5-110B-chat Bai et al. (2023) 02/24 1164 80.2 100 B Qianwen LICENSE
China Alibaba qwen1.5-72B-chat Bai et al. (2023) 02/24 1152 77.4 72 B Qianwen LICENSE
China Alibaba qwen1.5-32B-chat Bai et al. (2023) 02/24 1126 74.3 32 B Qianwen LICENSE
@@
Table 4: Additional council variations consisting of a hand-picked subset of LLMs. Flagships: the largest LLM from each organization; Smalls: the smallest LLMs from each organization; and Top-4, the top 4 LLMs according to Chatbot Arena as of May 2024.
Appendix B Additional Findings
all 0.92 (+0.01) 0.05 (+0.04) 1.0 (+0.48) 0.81 (+0.27) 0.35 (-0.19)
top-4 0.86 (+0.07) 0.04 (+0.03) 1.0 (+0.48) 0.84 (+0.24) 0.25 (-0.13)
Table 5: Key judging qualities from councils (no aggregation) for hand-selected sub-councils (Table 3) with only positionally consistent votes (positionally inconsistent ratings are filtered out prior to ranking). The value in parentheses represents the change compared to using all votes without a first filtering step.
Can we proactively remove inconsistent votes?
Table 1 shows a wide variety of judging qualities within the council, especially in consistency and separability. In the two-game setup, we gather ratings for models in both positions during pairwise comparisons, allowing us to identify and potentially remove inconsistent ratings before calculating ELO scores.
Automatically removing inconsistent votes affects the weighting of LLM judges, as those with more inconsistent ratings will have fewer votes counted and thus less influence overall. However, this could be a positive outcome, as positional consistency is often a sign that the judgment was arbitrary, especially in cases where tie ratings are not permitted. By comparison, some arena-based evaluation systems like Zheng et al. (2024) allow judges to declare a tie between responses, however these battles are completely excluded from ELO scoring.
In our emotional intelligence (EI) case study, we chose to retain all inconsistent votes, allowing downstream processes like aggregation, Bradley-Terry scoring, and bootstrapping to implement any diminished influence caused by inconsistency.
For the hand-selected sub-councils analyzed in Section 5.2, Table 5 shows the changes in key judging qualities when considering only consistent votes. Table 6 shows the impact of inconsistent vote pre-filtering on length bias.
Extended judging profiles and references to larger visualizations.
Due to table size constraints, we have divided the full tables of key judging qualities into several smaller tables. The full-sized 20x20 heatmaps from our LMC EI case study also span multiple pages. For ease of reference, all large tables and figures are compiled here:
  • Table 7 shows measures of bias for individual judges and the LMC as a whole.
  • Table 8 shows measures of agreement for individual LLM judges.
  • Table 9 shows polarization and affinity for individual LLM judges.
  • Figure 10 shows a heatmap of the affinities between LLM judges and LLM respondents.
  • Figure 11 shows a heatmap of the normalized affinities between LLM judges and LLM respondents (the LMC’s consensus affinity subtracted out).
  • Figure 12 shows a graph consisting of each LLM judge’s top 5 affinities.
  • Figure 13 shows a heatmap of Cohen’s $\kappa$ sidewise agreement scores.
  • Figure 14 shows a graph consisting of each LLM judge’s top 5 most agreeable other LLMs.
  • Figure 15 shows a heatmap of the estimated LLM vs. LLM win rates.
Models with $>$200 words All models
All votes Consistent votes All votes Consistent votes
Average judge 0.158 0.143 0.502 0.319
LMC (majority) 0.116 0.129 0.365 0.354
LMC (mean pool) 0.112 0.139 0.592 0.389
LMC (no aggregation) 0.125 0.106 0.545 0.347
@@ Appendix B Additional Findings
Figure 9: Comparison of full LMC rankings (no aggregation), with self-grading (left, default) permitted and with self-grading disabled (right). The rankings are broadly identical, confirming that the ensemble of LLM judges of the Language Model Council mitigates self-enhancement bias.
@@
Table 7: LMC judging profile relates for bias, with and without consistent votes.
@@
Table 8: LMC judging profiles related to agreement.
@@
Table 9: LMC judging profiles related to affinity.
Figure 10: Heatmap of the affinities of LLM judges to LLM respondents. The relatively consistent horizontal bands in the heatmap suggest a clear consensus on the preferred LLM participants. Some judges, like mistral-large, exhibit high polarization, with a significant difference between their highest and lowest-rated LLMs. In contrast, judges like llama3-8b display a narrower range of affinity. The highest affinity expressed by any LLM comes from mistral-large for qwen1.5-110B, while the strong horizontal blue band for gemini-1.5-pro indicates it was a consensus low performer.
Figure 11: This heatmap shows the normalized affinities of LLM judges toward LLM respondents, with the LMC’s consensus affinity subtracted out. Self-enhancement bias is now visible along the diagonal, where most LLMs exhibit some level of bias—though not all. Interestingly, six LLMs, including Claude-3-Opus and mistral-large, display negative self-enhancement bias, rating their own responses lower than the council’s consensus. For instance, mixtral-8x7b shows a strong preference for llama3-70b’s responses (+0.22 points above the consensus), but this affection is not reciprocated—llama3-70b actually rates mixtral-8x7b -0.08 points below the LMC’s consensus. Mistral-large is a particularly critical judge, rating 15 out of 20 LLMs more harshly than the LMC’s consensus. In contrast, Claude-3-Sonnet is much more favorable, expressing negative affinity for only one LLM, qwen-1.5-110B. The family blocks along the diagonal also reveal patterns of family-enhancing or self-deprecation bias. The llama3 family shows the highest family-enhancing bias, while the mistral family is more divided. Mistral-large and mistral-medium disproportionately rate their fellow family members, including themselves, more negatively, whereas mixtral-8x7b shows positive family bias.
Figure 12: Graph of top 5 affinities. An edge exists from LLM $a$ to LLM $b$ if $\text{affinity}(a,b)$ is in the top 5 affinities for LLM $a$. This view allows us to identify popular LLMs, hipster LLMs, and "LLM friendships" (where two LLMs have strong mutual affinity for each other). Popular LLMs, such as GPT-4o, have many arrows pointing toward them, while unpopular LLMs like dbrx-instruct have none. Qwen-1.5-32B and command-r form a "friendship" with mutual strong affinities, though Qwen-1.5-32B also receives incoming edges from three other LLMs, making it more widely liked. Some LLMs have only one "fan," others have several, and some have none. Of course, this analysis is somewhat contrived, as affinities are continuous values and using the top 5 as a cutoff is arbitrary. Nevertheless, it offers an interesting starting point for studying the dynamics and patterns that emerge when an arbitrary affinity threshold is established.
Figure 13: Heatmap of LLM judge Cohen’s $\kappa$ sidewise agreement scores. Since these are sidewise agreement scores, alignment on fine-grained ratings is not captured. However, the strong red band across the council rows indicates that the council is functioning as expected, representing a meaningful majority consensus. Inter-family agreement appears high overall, except within the Qwen-1.5 family, which shows lower scores of 0.03 and 0.08. In contrast, the OpenAI family demonstrates the highest inter-family agreement, with scores ranging from 0.30 to 0.49. Overall, LLM judges tend to be agreeing (no negative scores), though Llama-3-8b, Qwen-1.5-72b, and Gemini-1.0-Pro have the lowest agreement scores on the board. Interestingly, mistral-large and Claude-3-Opus show notably higher agreement scores than any other LLM pair.
Figure 14: Graph of top 5 agreement. An edge exists from LLM $a$ to LLM $b$ if $\text{agreement}(a,b)$ is in the top 5 agreement scores for LLM $a$. This visualization helps identify representative LLMs—those with many arrows pointing toward them are the ones that many other LLMs agree with. It also reveals communities of LLMs that tend to align with each other. While families exhibit high agreement in Figure 13, this graph shows fewer arrows within families, suggesting that certain non-family LLMs achieve higher agreement. A reverse graph, showing the bottom 5 agreement scores, could highlight contrarian LLMs. Overall, this approach helps identify the LLMs most agreed upon by other council members, which can be useful when selecting a representative sub-council.
Figure 15: Heatmap of the estimated LLM vs. LLM win rates. One notable outcome of using the Bradley-Terry Bradley and Terry (1952) estimation with a single reference model is the elimination of the "rock-paper-scissors" effect, where LLMs might have disproportionately favorable matchups. With a single reference model, the estimated win rates across all model pairs remain perfectly consistent, ensuring that while win rates vary, the distributional variance of these rates remains fixed. This design promotes stable relative rankings, which is desirable for evaluation purposes. However, it likely deviates from real-world scenarios where head-to-head win rates would be more heterogeneous, with different LLMs having specific dynamic advantages over others.
Appendix C LLM Judge Calibration
To understand the reliability and natural variability of LLM model judges, we collect pairwise preference ratings on three responses to the same interpersonal conflict using different temperatures and pairwise comparison options. Two responses are competitive, and one is intentionally generic to serve as a ranking baseline (Figure 17). We measure invariability and Pairwise Positional Consistency (PPC), defined below.
C.1 Invariability
How reliably does the model give the same preference with the same pair of responses in the same order?
Let:
  • $P$ be the set of all pairs of responses.
  • $n$ be the number of repetitions.
The frequency of the mode for the $i$-th pair is given by:
$f_{i}=\frac{\sum_{j=1}^{n}\mathbb{I}(R_{i,j}=\text{mode}(R_{i}))}{n}$
$invariability=\frac{1}{|P|}\sum_{i\in P}f_{i}$
C.2 Pairwise Positional Consistency (PPC)
How reliably does the model give a consistent preference with the same pair of responses in swapped order?
A rating couplet consists of a single rating of a pair of responses and then a rating of the same pair of responses in swapped order. For multiple repetitions of the same pair of responses in both orders, we take the percentage of consistent couplets over all possible rating couplets to factor out spuriously inconsistent couplets.
Let:
  • $P$ be the set of all pairs of responses.
  • $n$ be the number of repetitions.
[…]
$ppc=\frac{1}{|P|\cdot n^{2}}\sum_{i\in P}\sum_{j=1}^{n}\sum_{k=1}^{n}\text{are\_consistent}(R_{i,j},R_{i^{\prime},k})$
C.3 Experiment
Each LLM judge is prompted 5 times with the original pairwise comparison prompt (Figure 30) and 5 times with a trivially reworded version of the prompt. [^14] Trivial rewording involves changing the first sentence of the judging prompt (Figure 30) to: ”This person is experiencing an emotional dilemma and is seeking guidance and help.” This is repeated for the swapped order of responses.
For a single pair of responses, there are 10 reps (5 reps for each prompt * 2 prompts) in one order and 10 reps in the swapped order. Thus, there are $10*10=100$ possible rating couplets, which forms the denominator for PPC calculation.
The are_consistent function for consistency metrics is based on the mapping defined in Table 10.
A>>B or A>B A>>B or A>B FALSE TRUE TRUE FALSE
A>>B or A>B B>>A or B>A TRUE FALSE FALSE FALSE
A>>B or A>B Asimilar-$\sim$=B FALSE TRUE TRUE FALSE
B>>A or B>A A>>B or A>B TRUE FALSE FALSE FALSE
B>>A or B>A B>>A or B>A FALSE TRUE FALSE TRUE
B>>A or B>A Asimilar-$\sim$=B FALSE TRUE FALSE TRUE
Asimilar-$\sim$=B A>>B or A>B FALSE TRUE TRUE FALSE
Asimilar-$\sim$=B B>>A or B>A FALSE TRUE FALSE TRUE
Asimilar-$\sim$=B Asimilar-$\sim$=B TRUE FALSE FALSE FALSE
Table 10: Reference table for categorizing a couplet of order-swapped ratings of the same set of items, (A, B) vs. (B, A). Consistency is still counted as long as the overall side of the preference is consistent. Position-inconsistent ratings are either biased towards the first or second position.
  1. 1.
    Coarse preferences with tie option (A>B, B>A, Asimilar-$\sim$=B)
  2. 2.
    Coarse preferences without tie option (A>B, B>A)
  3. 3.
  4. 4.
    Granular preferences without tie option (A>>B, A>B, B>A, B>>A)
Figure 16: Calibration scores for invariability (left) and pairwise positional consistency (PPC) (right), averaged over 20 LLMs and 10 repetitions for each under different pairwise comparison options.
C.4 Results
temperature=0 is the best for reliability. The best average invariability and pairwise positional consistency across all 20 LLMs on the council is achieved with $temperature=0$. To our surprise, only $13/20$ models produce fully invariant ratings, even with $temperature=0$.
Coarse or granular rating options? Under $temperature=0$, the invariability between using coarse and granular rating options is small (0.98 vs. 0.96). The difference in PPC is more stark (0.91 vs. 0.80), though still tolerable. We decide to proceed with granular rating options for the main experiment for parity with Arena Hard Li et al. (2024a) to give more weight to strong preferences in the ELO calculation.
To include or not include a tie? Excluding the tie option slightly improves invariability and PPC at some temperatures, with negligible negative impact using $temperature=0$.
Figure 16 shows calibration scores for invariability and PPC, averaged over 20 LLMs and 10 repetitions for each under different pairwise comparison options. Table 11 shows a detailed breakdown per LLM, using granular pairwise comparison options without ties and with $temperature=0$.
C.5 Conclusion
Our calibration study concludes with the decision to use granular comparison options without a tie to "force" judges to choose a side, thereby better distinguishing models, and with temperature 0.
(strong votes) (first) (second)
claude-3-haiku 100.0% 50.0% 50.0% 50.0% 50.0%
claude-3-opus 100.0% 50.0% 100.0% 0.0% 0.0%
claude-3-sonnet 100.0% 50.0% 100.0% 0.0% 0.0%
command-r 100.0% 50.0% 50.0% 50.0% 50.0%
command-r-plus 100.0% 50.0% 100.0% 0.0% 0.0%
mistral-large 100.0% 50.0% 50.0% 0.0% 0.0%
mistral-medium 100.0% 50.0% 50.0% 0.0% 0.0%
mixtral-8x7b 100.0% 25.0% 50.0% 0.0% 0.0%
gpt-3.5-turbo-0125 82.5% 50.0% 95.0% 0.0% 0.0%
gpt-4-0613 100.0% 50.0% 100.0% 0.0% 0.0%
gpt-4-turbo-2024-04-09 92.5% 50.0% 100.0% 0.0% 0.0%
gpt-4o-2024-05-13 95.0% 50.0% 100.0% 0.0% 0.0%
qwen1.5-110B-Chat 100.0% 50.0% 100.0% 0.0% 0.0%
qwen1.5-32B-Chat 95.0% 25.0% 100.0% 0.0% 0.0%
qwen1.5-72B-Chat 100.0% 50.0% 50.0% 0.0% 0.0%
dbrx-instruct 92.5% 50.0% 65.0% 50.0% 50.0%
llama-3-70b-chat 100.0% 50.0% 100.0% 0.0% 0.0%
llama-3-8b-chat 82.5% 50.0% 50.0% 50.0% 50.0%
gemini-1.0-pro 75.0% 25.0% 69.5% 0.0% 0.0%
gemini-1.5-pro 100.0% 50.0% 100.0% 0.0% 0.0%
Table 11: Judging calibration results for 20 LLMs with using granular comparison options without a tie, with $temperature=0$. This is the same setting that was used for the paper’s primary case study (Section 3).
Appendix D Details on Reference Model Selection for the EI Case Study
Figure 18: Rankings from the dry run with 5% of the data with GPT-4-0613 (left) or Qwen-1.5-32B-Chat (right) as the reference model. When the reference model is uncompetitive, separability among top performing models is poor.
D.1 Understanding the Bradley-Terry Procedure
In a naive arena procedure with pairwise comparisons, every model’s response is paired with every other model’s response, requiring $O(n^{2})$ comparisons for $n$ models. This approach is resource-intensive and impractical for large $n$. To circumvent the need for a quadratic number of comparisons, the Bradley-Terry algorithm Bradley and Terry (1952) can be employed to determine expected win rates among a group of models, even without direct head-to-head battles between every pair.
The Bradley-Terry model offers a statistical method to estimate the relative strengths or “abilities” of items (models) based on pairwise comparison data. The key components of the model are:
  • Skill Parameters: Each model $i$ is assigned a positive real-valued parameter $\pi_{i}$, representing its skill or ability.
  • Win Probability: The probability that model $i$ beats model $j$ is given by:
    $P(i\text{ beats }j)=\frac{\pi_{i}}{\pi_{i}+\pi_{j}}$ (1)
Estimating Skill Parameters with Incomplete Data
Even without direct comparisons between every pair of models, we can estimate the skill parameters $\{\pi_{i}\}$ using available comparison data through the following steps:
  1. 1.
    Collect Pairwise Comparisons: Perform a subset of all possible pairwise comparisons, resulting in observed outcomes (which model won against which).
  2. 2.
    Set Up Likelihood Equations: The likelihood of the observed data, given the skill parameters, is formulated based on the Bradley-Terry probabilities.
  3. 3.
    Maximum Likelihood Estimation (MLE):
    • Objective: Find the set of skill parameters $\{\pi_{i}\}$ that maximize the likelihood of the observed data.
    • Process: Solve the likelihood equations derived from the comparisons to estimate the $\{\pi_{i}\}$.
  4. 4.
    Compute Expected Win Rates:
    • With the estimated skill parameters, calculate the expected probability that any model $i$ beats any model $j$ using:
      $P(i\text{ beats }j)=\frac{\pi_{i}}{\pi_{i}+\pi_{j}}$ (2)
    • Note: This computation is valid even for pairs of models that were not directly compared.
The Bradley-Terry algorithm enables us to estimate the expected win rates between all pairs of models without requiring a prohibitive number of direct comparisons by:
  • Assigning a skill parameter to each model.
  • Using observed pairwise comparisons to estimate these parameters via maximum likelihood estimation.
  • Calculating the probabilities of any model defeating another using the estimated parameters.
D.2 What is the reference model?
The reference model is the model whose responses are used across all pairwise comparisons. In this way, the reference model serves as a shared anchor for all models to be evaluated against. For example, if we have models W, X, Y, and Z, and use model Z as the reference, the pairwise comparisons are: (W vs. Z), (X vs. Z), and (Y vs. Z). The use of a reference model requires $O(n)$ comparisons for $n$ models.
While the Bradley-Terry algorithm doesn’t require a shared reference model, using one ensures a more consistent win rate estimation across all model pairs.
D.3 Selecting our reference model
In our arena-based LMC EI case study, we use a single reference model, following the approach of other arena-based benchmarks like Chatbot Arena Hard Li et al. (2024a) and Alpaca Eval Dubois et al. (2024a), which used GPT-4-0314 and GPT-4-turbo, respectively. It is unclear how Li et al. (2024a) and Dubois et al. (2024a) chose their reference models, but we will explain our choice.
We initially used GPT-4-0613 from OpenAI as our reference model. In a dry run, we observed poor separability in ELO scores (Figure 18). Other models won very often against GPT-4-0613, (we believe this was due to its average response length of 173 words, which was much less than the suggested 250-word limit (Table 1)). When the reference model is uncompetitive, ELO scores for other models get inflated, reducing ranking separability.
This led us to a key realization that for better separability, the reference model should have a varied mix of wins and losses against other models. We chose Qwen-1.5-32B as an alternative because it ranked mid-range in the initial dry run (Figure 18). Redoing the dry run with Qwen-1.5-32B improved separability substantially, so we proceeded to use it for the main experiment.
A more systematic approach to selecting the reference model—like having more randomized matchups, or using multiple reference models—could strengthen our arena-based LLM evaluation methods. However, this is beyond our paper’s scope and budget.
The choice of Qwen-1.5-32B as our reference model is not cherrypicking (the opposite actually). We believe that the risk of invalidating our key findings due to this choice is low. However, we speculate in our main Results that choosing the smaller Qwen model may have given an outsized advantage to larger models in the same family (Qwen-1.5-110B in particular).
Appendix E Human Evaluation